OpenStudy (destro):

Use a transformation to solve the equation. z – (–0.3) = –2.4 A. z = –2.7 B. z = –2.1 C. z = 2.1 D. z = 2.7

OpenStudy (shinalcantara):

z - (-0.3) = -2.4 -(-) is + z + 0.3 = -2.4 transpose 0.3 to the right side of the equation z = -2.4 - 0.3 z = ?